### 15 Feb 2024 v1

# An Evaluation of Real-time Adaptive Sampling Change Point Detection Algorithm using KCUSUM

Abstract.
Detecting abrupt changes in real-time data streams from scientific simulations presents a challenging task, demanding the deployment of accurate and efficient algorithms. Identifying change points in live data stream involves continuous scrutiny of incoming observations for deviations in their statistical characteristics, particularly in high-volume data scenarios. Maintaining a balance between sudden change detection and minimizing false alarms is vital. Many existing algorithms for this purpose rely on known probability distributions, limiting their feasibility. In this study, we introduce the Kernel-based Cumulative Sum (KCUSUM) algorithm, a non-parametric extension of the traditional Cumulative Sum (CUSUM) method, which has gained prominence for its efficacy in online change point detection under less restrictive conditions. KCUSUM splits itself by comparing incoming samples directly with reference samples and computes a statistic grounded in the Maximum Mean Discrepancy (MMD) non-parametric framework. This approach extends KCUSUM’s pertinence to scenarios where only reference samples are available, such as atomic trajectories of proteins in vacuum, facilitating the detection of deviations from the reference sample without prior knowledge of the data’s underlying distribution. Furthermore, by harnessing MMD’s inherent random-walk structure, we can theoretically analyze KCUSUM’s performance across various use cases, including metrics like expected delay and mean runtime to false alarms. Finally, we discuss real-world use cases from scientific simulations such as NWChem CODAR and protein folding data, demonstrating KCUSUM’s practical effectiveness in online change point detection.
Change point detection, Adaptive sampling, Molecular Dynamics (MD), Streaming data
ccs: Theory of computation Design and analysis of algorithmsccs: Theory of computation Online learning algorithms
1. Introduction
The detection of change points on streaming data without distributional information is an ongoing effort, with further difficulty introduced by increasingly larger volumes of data. In our work, we are interested in mitigating the potential issues imposed by these conditions, while still maintaining robust change detection. Detecting change points on live data could mean detecting a change from the average value or variance, or any other statistical property. In particular, the intended application on streaming data requires a change to be detected shortly after it occurs, without knowledge of the entire dataset. Examples of continuous datasets include quality control systems, scientific equipment, or biochemical simulations, among others.
There may be cases where sufficient knowledge of the data is available, for which an optimal algorithm can be chosen for live change detection. For example, if the probability distributions of before and after the change are known, the Cumulative Sum (CUSUM) algorithm (Algorithm II.I) is an optimal procedure, due to its short delay and minimal false alarm rate (Lorden, 1971, 1970). Apart from optimality, CUSUM is a simple algorithm to program and its output of maximum likelihood is easy to interpret. In some cases, the relevant statistical properties cannot be modeled directly from the data, and would be difficult to implement CUSUM without significant modification.
A similar topic to change detection is statistical hypothesis testing. In our work, we leverage established methods for two-sample testing and adjust them for change detection. In statistics, two-sample tests are non-parametric evaluations used to determine whether two datasets come from the same distribution. For the task of online change detection, our approach relies on kernel embedding to calculate the Maximum Mean Discrepancy (MMD) statistic. By using the MMD framework, two datasets can be compared by computing the distance between corresponding observable, induced by a positive definite kernel, which is introduced in Section 3. The distance found with MMD is an unbiased estimator and is not restricted to Euclidean datasets. This framework can be further generalized to hypothesis testing that involves strings, graphs or other structured datatypes (Muandet et al., [n. d.]).
Motivated by previous work using the MMD framework in hypothesis testing, we introduce our Kernel Cumulative Sum (KCUSUM) algorithm (Algorithm V.I). In contrast to CUSUM, KCUSUM does not require knowledge of the sample distribution before and after a change occurs. Instead, KCUSUM relies on samples from the pre-change regime and continuously compares them to incoming data using a kernel chosen by the user. By utilizing this framework, KCUSUM can detect changes to any distribution when the distance from the pre-change distribution exceeds a user-defined threshold. Initial theoretical results concern detection delay and average time to false alarm of the proposed algorithm. These results show the derived upper bound for detection time and the expected time until false alarm when no changes are present. Our analysis builds on existing theory for the CUSUM algorithm (Lorden, 1971, 1970).
The rest of our work is presented as follows. In the next section, we review some basic properties and behaviors of the CUSUM algorithm. In Section 3, the MMD framework is introduced. KCUSUM is reviewed in Section 5 where further analysis is presented. In Section 6, we show the results of using the KCUSUM algorithm on several datasets that cover expected use cases.
2. Cumulative Sum Algorithm
Let us observe a sequence of random variables $X_{n\geq 1}$ with index t such that for all 1 $\leq i<t$, the random variables $X_{i}$ have the distribution $p_{o}$, and for $i\geq t$ have the distribution $p_{1}$. We assume these variables have values in Euclidean space $\mathbb{R}^{d}$, although the proposed kernel variant is not restricted to these parameters. The index at $t$ is the change point with this convention, which is the detection target when performing analysis with the CUSUM algorithm. The analysis is performed in real time, where CUSUM calculates if a change has occurred by time $n$, on data available up to time $n$.
The Cumulative Sum algorithm, or CUSUM, is a live change point detection method proposed by Page (Page, 1954). To properly introduce CUSUM, we assume our two distributions, $p_{o}$ and $p_{1}$, have the densities $f_{o}$ and $f_{1}$, respectively. The full procedure of CUSUM is shown in ALGO II.I. At step $n$, a data point $X_{n}$ is observed, and the log-likelihood ratio $log\frac{f_{o}(x)}{f_{1}(x)}$ is calculated, with the result added onto the statistic $Z_{n}$. If the result is negative, then $Z_{n}$ is set to zero, effectively restarting the algorithm. If Zn crosses a threshold $h$, then a change is declared at time $n$. Formally, the CUSUM stopping time is:
$T_{CUSUM}=inf(n\geq 1|Z_{n}\geq h)$
To better see how the CUSUM algorithm works, consider the behavior of the log likelihood ratio before and after the change point t. Before the change, it has mean $-dKL(p_{o},p_{1})$, where $dKL(p_{o},p_{1})=\mathbb{E}_{P_{o}}log\frac{f_{o}(x)}{f_{1}(x)}$ is the Kullback-Liebler divergence between the two distributions. Since $dKL$ is positive when $p_{o}\neq p_{1}$, the increment term will have a negative average. This negative drift combined with the barrier at zero causes the statistic to remain near zero before any change occurs. After the change, the increment has a positive mean equal to $dKL(po,p1)$ and $Z_{n}$ begins to increase, eventually crossing any positive threshold h with probability one, ending the algorithm. Beyond these heuristic arguments, the CUSUM algorithm can be shown to be optimal in a certain sense, as we review below.
Algorithm 1 -II.I: Cumulative Sum (CUSUM)
input: Data $x_{1},x_{2},\ldots$ and threshold $h\geq 0$
initialize $Z_{o}=0$
For $n=1,2,\ldots$do
    $Z_{n}=\max\{0,Z_{n-1}+log\frac{f_{1}(x_{n})}{f_{o}(x_{n})}\}$
     if $Z_{n}\geq h$ then set $T_{CUSUM}=n$ and exit
    else continue
end
Example  II.I
Consider a normal distribution $N(a,b)$ with mean a and variance $b$. Let us assume we are detecting a change in the variance of this distribution, with pre-change parameters of $N(1,1)$ and post-change parameters of $N(1,4)$. A sample sequence $X1,...,Xn$ of length $n=400$ with the change in variance at $t=200$ is generated for this example. The change time is marked by a dashed line. The log likelihood ratio in this case is $log\frac{f_{o}(x)}{f_{1}(x)}=\frac{3}{8}x^{2}-\frac{3}{4}x+log\frac{1}{2}+\frac{3}{8}$. It can be seen that this value has a negative mean before the change occurs, and a positive mean after $t=200$. The threshold for CUSUM, $Z_{n}$, gives a detection time at $T_{CUSUM}=212$ with $h=10$.
In this section, we review the performance characteristics of the CUSUM algorithm. Each possible change time defines a different distribution on the sequences $\{X_{n}\}_{n\geq 1}$. If no change occurs, the variables are independent but identically distributed (i.i.d.) with $X_{i}~{}p_{o}$ for all $i\geq 1$. This distribution is denoted on sequences by $\mathbb{P}_{\infty}$, with the respective expectations $\mathbb{E}_{\infty}$. Generally, a change occurring at $t\geq 1$ means that for all $1\leq i\leq t-1,Xi$ is i.i.d. with $X_{i}~{}p_{o}$, and for $i\geq t$, they are i.i.d. with $X_{i}~{}p_{1}$. Let us denote $\mathbb{P}_{t}$ as the probability distribution on sequences under the assumption of a change at time $t$, where $\mathbb{E}_{t}$ represents the expectation for this distribution. For $n\geq 1$, let $F_{n}$ be the $\theta$-algebra $F_{n}=\theta(X_{1},\ldots,X_{n})$. $F_{n}$ represents the information contained in observations up to and including time $n$. Formally, an online change detector can be represented as a stopping time with respect to the filtration $\{F_{n}\}_{n\geq 1}$, that we denote as $T$, which can be interpreted as the estimate for the change point.
When running a change detector on a particular sequence, two types of errors are possible. There may be a false alarm, which means the change is detected too early, or there may be a delay, meaning the change was detected late. We can formalize the levels of false alarm and delay by using the metrics of average run length to false alarm (ARL2FA) and worst case average detection delay (ESADD). These are standard metrics for evaluating change detectors (Lorden, 1971; Moustakides et al., 2009; Poor and Hadjiliadis, 2008).
For a change detector T, the average time to false alarm is:
(1) $ARL2FA=\mathbb{E}_{\infty}[T]$
ARL2FA is the average detection time given a sequence of observations without change. We can measure delay using Lorden’s criterion (Lorden, 1971). If $T$ is a change detector and there is a change at time $t\geq 1$, then the expected delay given the history of the observations up to time $t-1$ is the random variable $\mathbb{E}_{t}[(T-t)^{+}|F_{t-1}]^{1}$ We can obtain the worst case delay for a change at time $t$ by taking the essential supremum over all sequences of length $t-1$, denoted by esssup $\mathbb{E}_{t}[(T-t)+|F_{t-1}]$. By taking the supremum over all change times $t$, we find the worst case delay:
(2) $ESADD=sup_{1\leq t<\infty}\-\ esssup\-\ \mathbb{E}_{t}[(T-t)^{+}|\mathcal{F}_{t-1}]$
In particular, the CUSUM algorithm provides an optimal trade- off between the time to false alarm (ARL2FA) and the worst case delay (ESADD). This was first proved in an asymptotic form in (Lorden, 1971), and the result was later proved in full non-asymptotic form in (Moustakides, 1986). A proof of optimality can be found in (Moustakides, 1986; Poor and Hadjiliadis, 2008). Further information on the derivation and properties of the CUSUM may be found in (Basseville and Nikiforov, 1993), or (Poor and Hadjiliadis, 2008).
The close relation between the defined threshold $h$ and the performance levels ARL2FA and ESADD is non-trivial and involves solving numerically intractable integrals (Page, 1954; Moustakides, 1986). However, it is possible to derive some upper and lower bounds that may be useful for applications. For the sake of comparison with the analysis of KCUSUM, let us consider the following quantitative bounds on the performance of the CUSUM.
Proposition  II.II
The performance of the CUSUM algorithm can be shown to have the following bounds. If the average time to false alarm satisfies the inequality $ARL2FA_{CUSUM}\geq e^{h}$, then it can also be shown that $\mathbb{E}_{1}[((log\frac{f_{o}(x)}{f_{1}(x)})^{+})^{2}]<\infty$. Therefore, the following inequality is also true:
(3) $ESADD_{CUSUM}\leq\frac{h}{d_{KL}(p_{1},p_{o})}+\frac{h}{d_{KL}(p_{1},p_{o})^{2}}\mathbb{E}_{1}[((log\frac{f_{1}(x)}{f_{o}(x)})^{+})^{2}]$
One interpretation of these equations is that by increasing the user defined threshold $h$, we can increase the time until a false alarm at the cost of a longer detection delay. From the ESADD formula, we can see that the detection delay increases when the distributions get closer. The positive term of the log-likelihood ratio is related to the variance of the CUSUM statistic. In Corollary IV.III we shall obtain the somewhat analogous bounds for the KCUSUM algorithm. In the following section, we discuss the Maximum Mean Discrepancy (MMD) statistic before introducing the kernel variant CUSUM algorithm.
3. Maximum Mean Discrepancy
In statistics, a two-sample test is used to determine whether two datasets are drawn from the same distribution. One method to solve this problem is to consider the empirical measures within the datasets, and compute the distance between measures using probability metrics. If enough data is available, the empirical distance should approximate the true distance. As this distance becomes larger, we become more confident that the datasets come from distinct distributions. This concept is a basic principle underlying several classical testing methods, such as the Kolmogorov-Darling (Feller, 1948), Cramer-von-Mises (Anderson, 1962) and the Anderson-Darling (Pettitt, 1976) tests.
Similarly, our proposed MMD approach is also based on computing empirical distances between distributions. With MMD, the data is implicitly embedded in a Reproducing Kernel Hilbert Space (RKHS), that corresponds to a user-defined kernel function $k$ where the embedding distance is calculated (Gretton et al., 2012). Compared to classical testing methods, MMD leverages several features that makes it appealing for non-parametric evaluations. First, the computed MMD value has a range of unbiased estimators (SEE DEFINTION OF $P_{L}$). Secondly, additional versatility is offered by which kernel the user choses, which can applied to data without a natural Euclidean form, such as graphs, strings or other structured datatypes (Gärtner, 2003; Vishwanathan et al., 2008).
Assume that $\chi$ is a set, such that $k:\chi\times\chi\rightarrow\mathbb{R}$ is a kernel on this set. This kernel is a symmetric, positive definite function, that can be regarded as a measure of similarity. One may consider the set $\chi=\mathbb{R}^{n}$ and the Gaussian kernel:
(4) $k(x,y)=e^{\frac{-||x-y||^{2}}{2}}$
Other choices for the kernel function are shown in (Gretton et al., 2012). Let us assume that $\chi$ has the structure of measurable space $(\chi,\Sigma)$ and that the chosen kernel $k$ is a measurable function on $\chi\times\chi$ with the $\theta$-algebra product. We define $P(\chi)$ to be the set of all probability measures on $(\chi,\Sigma)$, and by using $k$, the subset $P_{k}=\{\mu\in P(\chi)|\mathbb{E}_{\chi~{}\mu}[\sqrt{k(x,x)}]<\infty$ is defined. If the chosen kernel is characteristic, then we can define the MMD statistic on $P_{k}(\chi)$, given as $d_{k}$. This value is given by the formula:
(5) $d_{k}(p_{o},p_{1})=\sqrt{\mathbb{E}_{p_{o}xp_{o}}[k(x,x`)]+\mathbb{E}_{p_{1}xp_{1}}[k(y,y`)]-2\mathbb{E}_{p_{o}xp_{1}}[k(x,y)]}$
See (Gretton et al., 2012) for complete details
One of the unbiased estimators of $d^{2}k$ shown in (Gretton et al., 2012) is the linear statistic $p_{L}$ defined below. For convenience, the estimator is defined as the following function $g$:
(6) $g((x_{o},x_{1}),(y_{o},y_{1})=k(x_{o},x_{1})+k(y_{o},y_{1})-k(x_{o},y_{1})-k(x_{1},y_{o})$
Now, consider two datasets $X=\{x_{1},\ldots,x_{n}\}\subseteq\chi$ and $Y=\{y_{1},\ldots,y_{n}\}\subseteq\chi$. Then, $P_{L}(X,Y)$ becomes:
(7) $P_{L}(X,Y)=\frac{1}{n/2}\sum_{i=1}^{n/2}g((x_{2i-1},x_{2i}),(y_{2i-1},y_{2i}))$
One interesting behavior to note, is that the linear statistic is the sum of i.i.d. terms, which means the central limit theorem may be used to approximate the distribution, which can inform tuning of the thresholds when MMD is used for hypothesis testing (Gretton et al., 2012). Furthermore, in the online setting, we can study the trajectory of the statistic (that is, as a function of $n$) using the theory of sums of i.i.d. random variables, greatly facilitating analysis. We will exploit this structure in the analysis of KCUSUM in the following section.
4. Related Work
The topic of online change point detection has its roots in industrial quality control systems (Shewhart and Deming, 1939) and sequential analysis (Wald, 1947). In particular, the CUSUM procedure is closely related to the sequential probability ratio test (SPRT), which is a foundational online hypothesis testing algorithm (Wald, 1947). In contrast to online change detection methods, offline detection procedures must wait until the entire sequence is observed, and all data becomes available to make a decision. Kernel offline change point analysis was explored in (Harchaoui et al., 2008). Each theoretical change time is used to partition the dataset into two groups, consisting of prior observations and later observations, and the two resulting datasets are compared using a kernel based test. This is repeated for each possible change time. If one of the comparisons yields a significant discrepancy between the pre- and post- observation datasets, then a change is declared at the point where the largest difference was observed. This procedure is of interest as relies on distributional knowledge, basing its decisions on comparisons between disjoint sets of observations.
Several approaches to non-parametric online change detection have been proposed in (Brodsky and Darkhovsky, 1993, 2010). As shown in their work, they apply only to the case of detecting changes in mean. Another method of applying a kernel non-parametric test to online change detection is the sliding window approach, as pursued in (Li et al., 2015). In this method, at each time a decision regarding the change is made based on the distance between the most recent fixed-size block of data and a block from the reference distribution, using an MMD distance $d_{k}$. This can be seen as the kernel-based generalization of Shewhart control charts (Shewhart and Deming, 1939).
Additionally, current research (Polunchenko and Tartakovsky, 2010) aims to deploy AI-based change detectors on streaming data. It is argued that the CUSUM algorithm and its generalizations are simple neural networks, which can be leveraged with machine learning procedures for more robust change detection. By utilizing the CUSUM algorithm, this method attaches itself to rely on prior knowledge of the observed distribution. This limits practicality in our target use cases, where distributional information may be unavailable.
Besides CUSUM, there are other algorithms with optimality properties, most notably the Shiryaev-Roberts (SR) change detector (Fuh, 2003). While the CUSUM algorithm minimizes the worst case delay (ESADD), the SR detector minimizes a form of average delay. Like CUSUM, the SR test statistic takes a simple recursive form, so it may be possible to extend our algorithmic construction to SR-type detectors as well, but this is outside the scope of this paper.
5. Kernel Cumulative Sum Algorithm
The kernel variant of the Cumulative Sum algorithm (KCUSUM) blends features from the CUSUM procedure with the proposed MMD framework. Instead of taking each new observation $x_{n}$ to compute the log likelihood ratio, which estimates the KL-divergence $dKL$, we use the new observations and reference samples to calculate an estimate of the MMD distance $d_{k}$.
We can define the KCUSUM algorithm with the help of our previously defined MMD statistic. Given $\delta>0$ ($\delta$ is detailed below), we can define $g_{\delta}$ as:
(8) $g_{\delta}((x_{o},x_{1}),(y_{o},y_{1}))=k(x_{o},x_{1})+k(y_{o},y_{1})-k(x_{o},y_{1})-k(x_{1},y_{o})-\delta$
The full step procedure of the KCUSUM algorithm is listed in Algorithm V.1. We assume that the change to be detected is from a reference distribution $p_{o}$ to an unknown distribution $p_{o}$. At even numbered iterations, the current observation $x_{n}$ is paired with the previous observation $x_{n-1}$. Then, the two points are compared with two reference samples $y_{n},y_{n-1}$ with MMD. The constant $\delta>0$ is subtracted from the result to get $v_{n}$. The variable $v_{n}$ is then added to the threshold statistic $Z_{n-1}$ to get the next value of $Z_{n}$. If the new value $Z_{n}$ would be negative, the algorithm resets to 0, effectively restarting the procedure. At odd numbered iterations, the $Z_{n}$ is unchanged. Note that as a consequence, KCUSUM can only raise an alarm at even numbered iterations.
The motivation behind subtracting a positive amount $\delta$ at each step of Algorithm IV.I is to guarantee that the increments of $v_{n}$ have a negative drift under the pre-change regime and positive drift in the post-change regime. This is a consistency property that enables us to formulate non-trivial upper and lower bounds on performance. Using the definitions in Algorithm IV.1, it is evident that before a change, $\mathbb{E}[v_{n}]=-\delta<0$ and after a change, $\mathbb{E}[v_{n}]=d^{2}k(p_{o},p_{1})-\delta$. Theoretically, the algorithm can detect a change to any distribution $p_{1}$ that is at least distance $\surd\delta$ away from the reference distribution $p_{o}$.
Example  V.I
Let us consider detecting a change in variance of normally distributed variables, as in Example II.1. We have again the incoming stream of observations and our reference sample, with a change at $t=200$ shown as the red dashed line. The KCUSUM algorithm is based on the linear statistic $P_{L}$, so at each time $n$, the MMD estimate $k(x_{n-1},x_{n})+k(y_{n-1},y_{n})-k(x_{n},y_{n-1})-k(x_{n-1},y_{n})$ is computed. As in the CUSUM, a simple threshold $h$ is used to decide if a change has occurred. For this particular realization of the variables, a threshold of $h=5$ results in detection at $T_{KCUSUM}=225$. The value of $\delta$ was $\delta=\frac{1}{40}$.
Algorithm 2 -V.I: Kernel CUSUM (KCUSUM)
input:Thresholds $h\geq 0,\delta\geq$ and data $x_{1},x_{2},\ldots$
initialize $Z_{1}=0$
For $n=2,3,\ldots$do
    sample $y_{n}$ from reference measure $p_{o}$
     if $n$ is even then
      $v_{n}=g_{\delta}((x_{o},x_{1}),(y_{o},y_{1}))$
    Else
      $v_{n}$
    end
    $Z_{n}=max\{0,Z_{n-1}+v_{n}\}$
    if $Z_{n}\geq h$,then set $T_{KCUSUM}=n$ and exit
    Else continue
end
As found for the CUSUM algorithm, we now consider the rate of false alarms (ARL2FA) and average detection delay (ESADD). Using the filtration $\{F_{n}\}$ where $F_{n}=\sigma(x_{1},y_{1},\ldots,x_{n},y_{n})$, we can apply the previously defined Equations (1) and (2) to find ARL2FA and ESADD for the KCUSUM algorithm. To prove our upper and lower bounds, we adapt the methods of [1], which lets one reduce the problem of CUSUM analysis to the analysis of a random walk with i.i.d. terms. For convenience, we will group the variables together as $\{Z_{n}\}_{n\geq 1}$ where for $n\geq 1$:
(9) $Z_{n}=(x_{2n-1},x_{2n},y_{2n-1},y_{2n})$
This grouping reflects how KCUSUM processes data in pairs, and as a consequence, the bounds shown below involve additional factors of two compared to our CUSUM analysis (SEE PROPOSITION II.2). We define the associated auxiliary stopping times $c_{1},\ldots,c_{n}$ as:
(10) $c_{n}=inf\{k\geq n|\sum_{i=n}^{k}g_{\delta}(Z_{i})>h\}$
Theorem  V.II
Let $T_{KCUSUM}$ be the change detector corresponding to the described KCUSUM procedure (Algorithm IV.1). Let $p_{o}$ be the pre-change, or reference distribution. Then by using a threshold $\delta>0$, the time to false alarm is at least:
(11) $ARL2FA_{KCUSUM}\geq\frac{2}{\mathbb{P}_{\infty}(c_{1}<\infty)}$
If $p_{1}$ is a distribution with $d_{k}(p_{o},p_{1})>\surd\delta$, then the worst case detection delay is at most:
(12) $ESADD_{KCUSUM}\leq 2\mathbb{E}_{1}[c_{1}]$
Proof for Theorem  V.II
Let $b_{o}=0$ and for $n\geq 1$ let $b_{n}=max\{0,b_{n-1}+g_{\delta}(Z_{n})\}$. Define the stopping time $c$ as:
(13) $c=inf\{n\geq 1|b_{n}>h\}$
The relation between $c$ and the KCUSUM stopping time $T_{KCUSUM}$ is given as:
(14) $T_{KCUSUM}=2c$
Note, as discussed in [4], c can be represented as follows:
(15) $c=inf_{n\geq 1}c_{n}$
where each stopping time $c_{n}$ uses the same decision rule, and the only difference being that they operate on shifted versions of the input sequence $\{x_{i}\}i\geq 1$. In this way, Theorem 2 from [1] can be applied, which yields a lower bound of $c$ under $P_{\infty}$:
(16) $\mathbb{E}_{\infty}[c]\geq\frac{1}{\mathbb{P}_{\infty}(c_{1}<\infty)}$
Combining Equations (14) and (16) yields the claim (11). Now we consider bounding the worst case delay. If the sequence $\{x_{i}\}_{i\geq 1}$ has a change at an odd valued time, say $t=2m-1$ for $m\geq 1$, then the sequence $\{Z_{n}\}_{n\geq 1}$ has a change at $t=m$. Explicitly:
(17) $Z_{i}~{}\begin{cases}p_{o}\times p_{o}\times p_{o}\times p_{o}&\text{for $1\leq i<m$}\\ p_{1}\times p_{1}\times p_{o}\times p_{o}&\text{for $m\leq i$}\\ \end{cases}$
Now, here we can reason as in Theorem 2 of (Lorden, 1971):
(18) $\begin{split}\mathbb{E}_{2m-1}&[(T-(2m-1))^{+}]|\mathcal{F}_{2(m-1)}]\\ \stackrel{{\scriptstyle A}}{{=}}&\mathbb{E}_{2m-1}[(2c-2m+1)^{+}|\mathcal{F}_{2(m-1)}]\\ \stackrel{{\scriptstyle B}}{{\leq}}&\mathbb{E}_{2m-1}[(2c_{m}-2m+1)^{+}|\mathcal{F}_{2(m-1)}]\\ \leq&2\mathbb{E}_{2m-1}[(c_{m}-m+1)^{+}|\mathcal{F}_{2(m-1)}]\\ \stackrel{{\scriptstyle C}}{{=}}&2\mathbb{E}_{2m-1}[(c_{m}-m+1)^{+}]\\ \stackrel{{\scriptstyle D}}{{=}}&2\mathbb{E}_{1}[(c_{1}-1+1)^{+}]\\ =&2\mathbb{E}_{1}[c_{1}]\end{split}$
Step A follows from Equation (14) and Step B follows since $c$ is the infimum of the $\{c_{n}\}_{n\geq 1}$. Step C follows from the independence of $c_{m}$ from $F_{2(m-1)}$, and finally Step D follows from the fact that the distribution $c_{m}-m$ under $t=2m-1$ is the same as the distribution of $c_{1}-1$ under $t=1$.
The situation is slightly more complex if the change occurs a time t that is even, say $t=2m$ for some $m\geq 1$. In this case, the grouped sequence $\{Z_{n}\}_{n\geq 1}$ does not experience an abrupt change, but instead there are three relevant distributions. Specifically:
(19) $Z_{i}~{}\begin{cases}p_{o}\times p_{o}\times p_{o}\times p_{o}&\text{for $1\leq i<m$}\\ p_{o}\times p_{1}\times p_{o}\times p_{o}&\text{for $i=m$}\\ p_{1}\times p_{1}\times p_{o}\times p_{o}&\text{for $m<i$}\end{cases}$
By using the same reasoning in Equation (18), we have:
(20) $\begin{split}\mathbb{E}_{2m}&[(T-2m)^{+}]|\mathcal{F}_{2m-1}]\\ =&\mathbb{E}_{2m}[(2c-2m)^{+}|\mathcal{F}_{2m-1}]\\ \leq&2\mathbb{E}_{2m}[(c_{m+1}-m)^{+}|\mathcal{F}_{2m-1}]\\ =&2\mathbb{E}_{2m}[(c_{m+1}-m)^{+}]\\ =&2\mathbb{E}_{2}[(c_{2}-1)^{+}]\\ =&2\mathbb{E}_{1}[(1+c_{1}-1)^{+}]\\ =&2\mathbb{E}_{1}[c_{1}]\end{split}$
If we combine Equations (18) and (20), we can see that for all $t\geq 1$:
(21) $\mathbb{E}_{t}[(T-t)^{+}|\mathcal{F}_{t-1}]\leq 2\mathbb{E}_{1}[c_{1}]$
Combining Equation (21) with the definition of worst case delay shown in Equation (2) yields the inequality shown in Equation (12). We can combine Theorem IV.2 with certain facts about random walks (Lemma A.2 and Proposition A.1) to get more specific bounds on the delays and false alarms, as shown in the following Corollary. In this corollary, we assume that the kernel $k$ is bounded by a constant $||k||_{\infty}$, and also assume that $\delta$ is bounded by $\delta<2||k||_{\infty}$. This is a necessary assumption when the kernel is bounded, since if $\delta\geq 2||k||_{\infty}$, then $g_{\delta}(z)\leq 0$ for all $z$, and it will not be possible to detect any changes.
Corollary  V.III
Let the assumptions of Theorem IV.2 hold. Further assume that the kernel $k$ is bounded by a constant $||k||_{\infty}$ and let $\delta<2||k||_{\infty}$. Then the time to false alarm satisfies:
(22) $ARL2FA_{KCUSUM}\geq 2e^{\frac{h}{4||k||_{\infty}}log(1+\frac{\delta}{4||k||_{\infty}})}$
If $p_{1}$ is a distribution with $d_{k}(p_{0},p_{1})>\surd\delta$, then the worst case detection delay has a maximum of:
(23) $ESADD_{KCUSUM}\leq\frac{2h}{d_{k}(p_{o},p_{1})^{2}-\delta}+\frac{8||k||^{2}_{\infty}}{(d_{k}(p_{o},p_{1})^{2}-\delta)^{2}}$
Proof for Corollary  V.III
To start, note that:
(24) $\begin{split}\mathbb{P}_{\infty}(c_{1}<\infty)=\mathbb{P}_{\infty}(inf\{k\geq 1|\sum^{k}_{i=1}g_{\delta}(z_{i})>h\}<\infty)\\ =\mathbb{P}_{\infty}(sup_{k\geq 1}\sum^{k}_{i=1}g_{\delta}(z_{i})>h)\end{split}$
To upper-bound the last term in this equation, we will apply Lemma A.2. Note that under our assumption that the kernel is bounded, the moment generating function $M(r)=\mathbb{E}_{\infty}[e^{rg_{/delta}(Z_{1})}]$ is guaranteed to be well defined for all $r\geq 0$. Therefore:
(25) $\mathbb{P}_{\infty}(sup_{k\geq 1}\sum^{k}_{i=1}g_{\delta}(z_{i})>h)\leq e^{-rh}$
where $r$ is any number satisfying $r>0$ and $M(r)\leq 1$. To identify such an $r$, start with a second order expansion of $M$:
[…]
Under the assumption that $\delta<2||k||_{\infty}$, it holds that $|g_{\delta}(z)|\leq 4||k||_{\infty}$ and:
[…]
If we minimize the with right hand side with respect to $r$, it yields:
(28) $r=\frac{1}{4||k||_{\infty}}log(1+\frac{\delta}{4||k||_{\infty}})$
The equations (11), (24) and (25) can be combined with the previous definition of $r$ (28), which yields the result of (22). For the delay, note that $\mathbb{E}_{1}[c_{1}]$ is the expectation time of a random walk crossing the upper boundary, under positive drift. Therefore, we can apply Proposition A.1, which shows:
(29) $\begin{split}\mathbb{E}_{1}[c_{1}]\leq\frac{h}{\mathbb{E}_{1}[g_{\delta}(z_{1})]}+\frac{\mathbb{E}_{1}[(g_{\delta}(z_{1})^{+})^{2}]}{\mathbb{E}_{1}[g_{\delta}(z_{1})]^{2}}\\[10.0pt] =\frac{h}{d_{k}(p_{o},p_{1})^{2}-\delta}+\frac{\mathbb{E}_{1}[(g_{\delta}(z_{1})^{+})^{2}]}{(d_{k}(p_{o},p_{1})^{2}-\delta)^{2}}\end{split}$
By combining (12) and (29) with the bound $g_{\delta}(z)^{+}\leq 2||k||_{\infty}$, we find the relation (23). As discussed in Theorem V.II, there is a logarithmic relationship between delay and time to false alarm. If we compute the smallest value of $h$ that guarantees an alert rate $x$ for each level of false alarm $x\in\{1,2,\ldots,10^{4}\}$, the delay can be calculated according to (23). The analysis was performed on a theoretical problem with the parameters $d_{k}(p_{o},p_{1})^{2}=\frac{1}{6},\delta=2^{-}5$ and $||k||_{\infty}=0.5$.
6. Results
In this section, we evaluate the KCUSUM algorithm on a variety of change detection tasks. Motivated by our prior theoretical analysis and the results of (Flynn and Yoo, 2020), we apply the KCUSUM algorithm to more complex data structures. These structures allow the bounding of general performance for each type of task, and are expected use cases for the algorithm. For each case, observations consist of vectors in $\mathbb{R}^{4}$ based in Cartesian coordinates, and grouped by time. The reference data is a set of observable sampled from the pre-change distribution and is unique for each task.
  1. (1)
    Change of mean. Atomic collisions at thermal equilibrium with one scaled component
  2. (2)
    Change of mean. 1H9T protein with partial structure deformation.
  3. (3)
    Change of mean and variance. 1FME peptide with fast folding properties ($\beta-\beta-\alpha$).
Figure 1. Distribution of reference (orange) and testing data (blue) for Case 1. Each change point falls in between the set of dashed lines.
Figure 2. A sample evaluation chart of the KCUSUM algorithm on data from Case 2.
Figure 3. 1H9T structure before and after the change was detected.
Figure 4. A sample evaluation chart of the KCUSUM algorithm on data from Case 3.
Figure 5. 1FME structure before, during and after the change was detected.
Figure 6. The performance of the KCUSUM algorithm on the proposed change detection tasks.
In each case, a Monte Carlo approach is used to obtain performance bounds on KCUSUM. To test delay, each dataset is masked with a normally distributed (0,1) noise layer to generate $n=10000$ unique sequences. The algorithm is then run and each delay time is recorded, to obtain an estimate for ESADD. The actual delay time $t$ is independent for each task. To obtain an estimate for ARL2FA, we observe cases where a false alarm is raised before the real change point occurs, where the time is taken from the first observable.
For each case, we see a logarithmic relationship between delay and time to false alarm, which can vary in scale. This scale is an indicator of the performance of KCUSUM for each kind of task. From Figures 1-4, the Case 1 and 3 change points are more easily detected, while Case 2 proves more difficult for the algorithm to detect. It is important to note for each case, that the data input to the algorithm is the incoming observables and samples drawn from the pre-change distribution; there is no information given about the post-change distribution. These results suggests that KCUSUM may be a promising change detector for problems where there is little information about the kind of change that occurs. Additionally, when a change point is detected, the algorithm is reset so that it continues to process incoming data. When the reset occurs, a new reference sample is taken from the post-change distribution, which becomes our new reference. After a significant conformational change, the original reference sample would likely increase ARL2FA if used past the detected change point. The rate at which ARL2FA changes from reference selection is a current focus of our research.
7. Conclusion
In conclusion, the evaluation of the KCUSUM algorithm on various change point detection tasks has provided valuable insights into its performance across different molecular dynamics trajectory structures. The algorithm was applied to complex data structures based on prior theoretical analysis and reference data unique to each task. Three specific cases were examined, involving changes in mean, mean and variance, and complex atomic structures. A Monte Carlo approach was employed to establish performance bounds on KCUSUM for each case. The evaluation included testing delay through the introduction of noise layers to generate unique sequences. The algorithm’s performance was assessed in terms of delay and time to false alarm, revealing a logarithmic relationship between the two, with varying scales indicative of the algorithm’s effectiveness for each task. Notably, the algorithm demonstrated effectiveness in cases involving changes in mean and atomic structure deformation, while facing challenges in detecting changes in mean and variance for a specific peptide with fast folding properties. The logarithmic relationship observed suggests that KCUSUM may be well-suited for change point detection in scenarios where little information is available about the nature of the change. Real data evaluation was also conducted, with visual representations provided for each case. These figures illustrated the distribution of reference and testing data, as well as sample evaluation charts showcasing the algorithm’s performance. In summary, the theoretical analysis and empirical evaluations collectively suggest that KCUSUM holds promise as a change detector, particularly in situations where limited information is available about the type of change occurring. Further applications and testing in real-world scenarios will contribute to a deeper understanding of its capabilities and potential use cases.
Acknowledgements
This material is based upon work supported by the Advanced Scientific Computing Research Program in the U.S Department of Energy, Office of Science, under FAIR (Funding for Accelerated, inclusive research) Award number DE-FOA-0002931. We are grateful for the support received from the SHI SRP-HPC ECP project funded by the U.S DOE, and for the excellent mentorship from the BNL scientists.
Appendix
Proposition II.I: Assume $\{a_{i}\}_{i\geq 1}$ are i.i.d real-value random variables such that $\mathbb{E}[a_{i}]=\mu>0$ and $\mathbb{E}[(a_{i}^{+})^{2}]<\infty$. Let us define $S_{n}$ and $T$ for $a\leq 0\leq b$ such that $S-n=\sum^{n}_{i=1}a_{i}$ and $T=inf\{n\geq 1|S_{n}\notin[a,b]\}$. Now:
$\mathbb{E}[T]\leq\frac{(1-\alpha)b+a\alpha}{\mu}+\frac{\mathbb{E}[(a_{1}^{+})^{2}]}{{\mu}^{2}}$
where $\alpha=\mathbb{P}(S_{T}<a)$.
Lemma II.II
Assume $\{a_{i}\}_{i\geq 1}$ is a sequence of i.i.d. real-value random variables such that for $n\geq 1$, we can define the partial sum as $S_{n}=\sum^{n}_{i=1}a_{i}$. Let $M(r)=\mathbb{E}[e^{ra_{1}}]$ be our moment generating function for $a_{i}$ and assume for some $q>0$, there is a $M(q)\leq 1$. Now, for any $h\geq 0$, we have:
$\mathbb{P}(sup_{n\geq 1}S_{n}>h)\leq e^{-qh}$
$ProofofLemmaII.II:$
Let us define $Z-n=e^{qS_{n}}$ for all $n\geq 1$. Now, we can show for $n\geq 1$:
$\begin{split}\mathbb{E}[Z_{n+1}|Z_{n}]=\mathbb{E}[\prod^{n+1}_{i=1}e^{qa_{i}}|Z_{n}]\\ =Z_{n}\mathbb{E}[e^{qa_{n+1}}]\\ =Z_{n}M(q)\leq Z_{n}\end{split}$
Then, for all such $n\geq 1$, it holds that $Z_{n}\geq 0$ and $\mathbb{E}[|E|]=M(q)^{n}\leq 1$. Since $Z_{n}$ is a non-negative supermartingale, then we have:
$\begin{split}\mathbb{P}(sup_{n\geq 1}S_{n}>h)=\mathbb{P}(sup_{n\geq 1}Z_{n}>e^{-qh})\\ \leq\mathbb{E}[e^{qa_{1}}]e^{-qh}\\ \leq e^{-qh}\end{split}$
The first part in the previous derivation follows from the monotonic function $x\rightharpoonup e^{x}$, where the second part is a result of Theorem 7.8 (Gallager, 1995). The final line is a result of our assumption on $q$.
$ProofofProposition\hskip 2.84526ptII.I$:
Let us define $S_{1},S_{2},\ldots$ as:
$S_{n}=\sum^{n}_{i=1}log\frac{f_{1}(x_{i})}{f_{o}(x_{i})}$
with $\alpha=\mathbb{P}_{\infty}(sup_{n\geq 1}S_{n}>h)$. Theorem 2 (Lorden, 1971) shows that CUSUM follows:
$ARL2FA_{CUSUM}\geq\frac{1}{\alpha}$
Since $\mathbb{E}_{\infty}[log\frac{f_{1}(x_{i})}{f_{o}(x_{i})}]=-d_{KL}(p_{o},p_{1})<0$, $S_{n}$ is a random walk with a negative drift. The moment generating function can be shown as:
$M(r)=\mathbb{E}[e^{rlog\frac{f_{1}(x_{i})}{f_{o}(x_{i})}}]=\mathbb{E}_{\infty}[(\frac{f_{1}(x_{i})}{f_{o}(x_{i})})^{r}]$
From $M(r)$, we can see that $g(1)=\mathbb{E}_{1}=1$. Now we can apply Lemma II.II with $q=1$ to show that $\alpha\leq e^{-h}$.
Now let $N$ be our stopping time $N=inf\{n\geq 1|S_{n}>h\}$. If we apply Theorem 2 (Lorden, 1971), it shows that:
$ESADD_{CUSUM}\leq\mathbb{E}_{1}[N]$
Assume that a change occurs at $t=1$, and that the variables $\frac{f_{1}(x_{i})}{f_{o}(x_{i})};i=1,2,\ldots$ are i.i.d. with a positive mean. Since $\{S_{n}\}_{n\geq 1}$ is a random walk with a positive drift $\mu=\mathbb{E}_{1}[log\frac{f_{1}(x_{i})}{f_{o}(x_{i})}]$, it follows from Proposition II.I with stopping time $N$ on the established bounds of ESADD.